# Thermal quasi-particle theory

### v1

@@ Abstract
The widely used thermal Hartree–Fock (HF) theory is generalized to include the effect of electron correlation while maintaining its quasi-independent-particle framework. An electron-correlated internal energy (or grand potential) is defined by the second-order finite-temperature many-body perturbation theory (MBPT), which then dictates the corresponding thermal orbital (quasi-particle) energies in such a way that all thermodynamic relations are obeyed. The associated density matrix is of the one-electron type, whose diagonal elements take the form of the Fermi–Dirac distribution functions, when the grand potential is minimized. The formulas for the entropy and chemical potential are unchanged from those of Fermi–Dirac or thermal HF theory. The theory thus postulates a finite-temperature extension of the second-order Dyson self-energy of one-particle many-body Green’s function theory and can be viewed as a second-order, diagonal, frequency-independent, thermal inverse Dyson equation. At low temperature, the theory approaches finite-temperature MBPT of the same order, but it outperforms the latter at intermediate temperature by including additional electron-correlation effects through orbital energies. A physical meaning of these thermal orbital energies (including that of thermal HF orbital energies, which has been elusive) is proposed.
@@ I Introduction
Thermal Hartree–Fock (HF) theoryHusimi (1940); Mermin (1963); Gu and Hirata (2024) is a curious ansatz. It uses an auxiliary one-electron Hamiltonian to define its one-electron density matrix, whereas a true density matrix for thermodynamics governed by an interacting Hamiltonian should be much higher ranked.Farid et al. (2000) Its diagonal elements, as variational parameters, ultimately become the Fermi–Dirac distribution functions of the non-interacting Fermi–Dirac theory, when the grand potential is minimized. The state energies defining this grand potential are, however, evaluated with the exact Hamiltonian containing two-electron interactions. Stipulated in this way, the theory does not seem to correspond to a single, consistent grand partition function.
@@
Nevertheless, the thermodynamic functions of this theory are shownArgyres et al. (1974); Gu and Hirata (2024) to obey all thermodynamic relations. They also correctly reduceGu and Hirata (2024) to the zero-temperature HF theory.Szabo and Ostlund (1982); Shavitt and Bartlett (2009) They are widely used in applications to, e.g., energy bands in a solid where their temperature-dependent orbital energies are often invoked, somewhat unscrupulously, in explaining metal-insulator transitions,Hermes and Hirata (2015) etc. This is in spite of the facts that their physical meaning is still obscurePain (2011) and that state energies in quantum mechanics are supposed to be constant of temperature.
@@
In this article, inspired by the immense success of thermal HF theory, we generalize this ansatz to include the effects of electron correlation, while keeping to its quasi-independent-particle framework and grand canonical ensemble. Our strategy of postulating what we call thermal quasi-particle theory is as follows:
@@
We first assume that every state consists of non-interacting quasi-particles, each occupying a well-defined orbital having an electron-correlated orbital energy in a Slater determinant. The internal energy (or equivalently, the grand potential) is postulated by consulting with finite-temperature many-body perturbation theory (MBPT).Hirata and Jha (2019, 2020); Hirata (2021a) Its definition dictates the forms of the thermal quasi-particle energies (i.e., the electron-correlated thermal orbital energies) such that all thermodynamic relations are obeyed. The corresponding density matrix is then of the one-electron type, whose diagonal elements take the form of the Fermi–Dirac distribution functions, when the grand potential is minimized. The formulas for the entropy and average number of electrons (determining the chemical potential) are unchanged from the corresponding formulas of Fermi–Dirac or thermal HF theory. The thermal quasi-particle energies thus defined are interpretable as a finite-temperature generalizationMatsubara (1955); Luttinger and Ward (1960); March et al. (1967); Kadanoff and Baym (2018) of the Dyson self-energies of one-particle many-body Green’s function (MBGF) theory.Hirata et al. (2017)
@@
Being based on perturbation theory, thermal quasi-particle theory forms a hierarchy of approximations with increasing accuracy and complexity. Fermi–Dirac and thermal HF theories can be viewed, respectively, as the zeroth- and first-order instances of this hierarchy. Here, we introduce a second-order thermal quasi-particle theory, which is the leading order in describing electron correlation, on the basis of the second-order grand potential or internal energy of finite-temperature MBPT.Hirata and Jha (2020); Hirata (2021a) We propose the expression for the corresponding second-order thermal self-energies, which obey all thermodynamic relations and ensure the variationality of the grand potential. The theory thus constitutes a second-order, diagonal, frequency-independent, thermal inverse Dyson equation.Hirata et al. (2017) It reduces to the second-order MBPT and MBGF (in the diagonal and frequency-independent approximation) at zero temperature.
@@
A comparison of thermal quasi-particle theories with finite-temperature MBPT suggests that the former may include higher-order perturbation corrections through correlation-corrected thermal orbital energies and can thereby outperform the latter at intermediate temperatures. Furthermore, a comparison with thermal full-configuration-interaction (FCI) theoryKou and Hirata (2014) supports a proposed physical meaning of these thermal orbital energies, including that of thermal HF orbital energies, which has been elusive.Pain (2011); Hirata (2021a)
II Fermi–Dirac theory
@@
Fermi–Dirac theory can be considered as the zeroth-order thermal quasi-particle theory [thermal QP(0) theory]. It is reviewed only briefly since it is discussed more fully in Ref. Gu and Hirata, 2024. Its internal energy $U^{(0)}$ and entropy $S^{(0)}$ at the inverse temperature $\beta=(k_{\text{B}}T)^{-1}$ are given by
@@
where $\bar{N}$ is the average number of electrons, and the summations are always taken over all spinorbitals. The $\langle E^{(n)}\rangle$ in Eq. (1) is a zeroth-order thermal average of state energies, i.e.,
@@
where $E_{I}^{(n)}$ is the $n$th-order perturbation correction to the energy of the $I$th Slater-determinant state according to Hirschfelder–Certain degenerate perturbation theoryHirschfelder and Certain (1974) with $E_{I}^{(0)}=\sum\epsilon_{i}^{(0)}$ ($i$ labels spinorbitals occupied in the $I$th state). Sum-over-states expressions of the thermal averages can be reduced to sum-over-orbital ones either by combinatorial logicHirata and Jha (2019, 2020) or by normal-ordered second quantization at finite temperature.Hirata (2021a)
@@
$\Omega^{(0)}$$\equiv$ $U^{(0)}-\mu^{(0)}\bar{N}-TS^{(0)}$ (8)
$=$ $\sum_{p}\left(\epsilon_{p}^{(0)}-\mu^{(0)}\right)f_{p}^{-}$
$+\frac{1}{\beta}\sum_{p}\left(f_{p}^{-}\ln f_{p}^{-}+f_{p}^{+}\ln f_{p}^{+}\right).$
As explicitly shown in Ref. Gu and Hirata, 2024, they satisfy the thermodynamic relations such as Eq. (8) and
@@
which is also easily verifiable.Gu and Hirata (2024)
@@
III Thermal Hartree–Fock theory
Thermal HF theoryHusimi (1940); Mermin (1963) is the first-order instance of thermal quasi-particle theory [thermal QP(1) theory], which is also fully discussed in Ref. Gu and Hirata, 2024. It accommodates the exact Hamiltonian with two-electron interactions within a quasi-independent-particle framework. Its internal energy $U^{\text{HF}}$ is postulated by an intuitively natural finite-temperature generalization of the zero-temperature HF energy,Hirata (2021a) i.e.,
@@ III Thermal Hartree–Fock theory
$\epsilon_{pq}^{\text{HF}}$ $=$ $h_{pq}+\sum_{r}\langle pr||qr\rangle f_{r}^{-}=\left\{\begin{array}[]{ll}\epsilon_{p}^{\text{HF}}&(p=q);\\ 0&(p\neq q),\end{array}\right.$ (17)
@@
The last identity proves that $\Omega^{\text{HF}}$ is indeed minimizedMermin (1963); Gu and Hirata (2024) by the orbitals that diagonalize $\bm{\epsilon}^{\text{HF}}$ [Eq. (17)] and by the one-particle density matrix whose eigenvalues are the Fermi–Dirac distribution functions of Eq. (18).
IV Thermal quasi-particle theory
Second-order thermal quasi-particle theory [thermal QP(2) theory] is postulated by its internal energy,
@@ IV Thermal quasi-particle theory
The last term $\langle E^{(2)}\rangle$ is the thermal average of second-order (degenerate) perturbation energiesHirschfelder and Certain (1974) and has been reduced in Eqs. (C7) and (C8) of Ref. Hirata and Jha, 2020 to the following sum-over-orbitals formula:
@@
It is a natural finite-temperature generalization of second-order many-body perturbation energy,Bartlett (1981); Shavitt and Bartlett (2009) but is considerably simpler than the full second-order correction to the internal energy $U^{(2)}$ of finite-temperature MBPT [Eq. (74) of Ref. Hirata and Jha, 2020]. The simplification is at least partly justified by the equally simpler treatments of the chemical-potential and entropy contributions to adhere to the quasi-independent-particle framework that the theory adopts. The first term of $\langle E^{(2)}\rangle$ [Eq. (LABEL:E2)] is the so-called non-HF term (or more precisely, non-thermal-HF term, in this case),Shavitt and Bartlett (2009) which is zero when thermal HF theory is used as the reference (where $F_{pq}=0$). The diagrammatic representationHirata (2021a) of $\langle E^{(2)}\rangle$ is given in Fig. 1.
@@
where $\mu^{\text{QP(2)}}$ is the chemical potential, which is determined by the same electroneutrality condition of Fermi–Dirac theory, i.e.,
@@
Here, $\Sigma_{pp}^{(2)}$ is a finite-temperature analogue of the second-order self-energy of MBGF.Hirata et al. (2017) In a zero-temperature reference, where $\epsilon_{p}^{(0)}$ is constant of $T$, it takes the form of
$\Sigma_{pp}^{(2)}$ $=$ $\sum_{q}^{\text{denom.}\neq 0}\frac{F_{qp}F_{pq}}{\epsilon_{p}^{(0)}-\epsilon_{q}^{(0)}}f_{q}^{+}-\sum_{q}^{\text{denom.}\neq 0}\frac{F_{qp}F_{pq}}{\epsilon_{q}^{(0)}-\epsilon_{p}^{(0)}}f_{q}^{-}$ (38)
$+\sum_{q,r}^{\text{denom.}\neq 0}\frac{\langle qr||pr\rangle F_{pq}}{\epsilon_{p}^{(0)}-\epsilon_{q}^{(0)}}f_{p}^{-}f_{q}^{+}$
$+\sum_{q,r}^{\text{denom.}\neq 0}\frac{F_{qp}\langle pr||qr\rangle}{\epsilon_{p}^{(0)}-\epsilon_{q}^{(0)}}f_{p}^{-}f_{q}^{+}$
@@
which is obtained by demanding all thermodynamic relations be obeyed by thermodynamic functions defined by this theory (see below for a derivation). A diagrammatic representation of $\Sigma_{pp}^{(2)}$ is given in Fig. 2.
@@
When a zero-temperature HF reference is used, in the $T=0$ limit, the first four terms vanish because $F_{pq}=0$, leaving only the last two terms as the zero-temperature second-order self-energy in the diagonal, frequency-independent approximation. Their diagrams (the last two in Fig. 2) are, therefore, isomorphic to the two second-order self-energy diagrams (e.g., Fig. 4 of Ref. Hirata et al., 2017), although their scope is different (the former are for $T\geq 0$, while the latter are only for $T=0$).
Figure 2: Second-order thermal Dyson self-energy diagrams in the zero-temperature reference [Eq. (38)], which are obtained by cutting an edge in each of the internal energy diagrams in Fig. 1 and trimming the dangling lines. See also the caption of Fig. 1 as well as Refs. Hirata, 2021a; Hirata et al., 2017.
@@
but not the explicit derivative $\partial f_{p}^{-}/\partial\mu^{\text{QP(2)}}$, which is a solution of a system of linear equations and cannot be written in a closed form.Gu and Hirata (2024) This justifies Eq. (34).
@@
which can be verified by using Eqs. (LABEL:E2) and (38). In fact, the form of $\Sigma_{pp}^{(2)}$ is determined in the first place from the general condition,
@@
which ensures that Eq. (43) and all similar relations be satisfied. The $\Sigma_{pp}^{(2)}$ formula of Eq. (38) was postulated by substituting Eq. (LABEL:E2) into the above formula.
Equation (44) is a finite-temperature generalizationMatsubara (1955); Luttinger and Ward (1960) of the $n$th-order self-energy of MBGF,Hirata et al. (2017, 2015) which is diagonal and frequency independent. Owing to the diagonal construction, the second- and higher-order thermal quasi-particle theory does not involve rotation of orbitals, unlike thermal HF theoryMermin (1963); Gu and Hirata (2024) or Feynman–Dyson MBGF,Hirata et al. (2017) which define the HF or Dyson orbitalsOrtiz (2020) as those that bring the Fock or self-energy matrix into a diagonal form. Despite the absence of orbital rotation, $\Omega^{\text{QP(2)}}$ is still minimized by varying $f_{p}^{-}$ (see below).
Diagrammatically, the differentiation with respect to $f_{p}^{-}$ corresponds to opening a closed, internal-energy diagram by deleting an edge.Hirata et al. (2017) A possible physical meaning of thermal self-energies or thermal quasi-particle energies, which are temperature dependent, is given and discussed in Sec. VI.
@@
$-k_{\text{B}}\beta^{2}\sum_{p}\frac{\partial\Sigma_{pp}^{(2)}}{\partial\beta}f_{p}^{-}+k_{\text{B}}\beta^{2}\frac{\partial\langle E^{(2)}\rangle}{\partial\beta}$
@@
where we used Eqs. (42) and (44), but not the explicit derivative $\partial f_{p}^{-}/\partial\beta$.Gu and Hirata (2024) This justifies Eq. (31).
@@
$\frac{\partial\Omega^{\text{QP(2)}}}{\partial f_{p}^{-}}$ $=$ $\left(\epsilon_{p}^{\text{HF}}-\mu^{\text{QP(2)}}\right)+\frac{1}{\beta}\ln\frac{f_{p}^{-}}{f_{p}^{+}}+\frac{\partial\langle E^{(2)}\rangle}{\partial f_{p}^{-}}$
@@
where Eq. (44) was used. This proves that the $f_{p}^{-}$ functions that minimize $\Omega^{\text{QP(2)}}$ take the forms of the Fermi–Dirac distribution functions, justifying Eqs. (35) and (36).
Insofar as Eq. (44) is satisfied (and thus all thermodynamic relations are obeyed and the variationality is ensured), there is considerable latitude in selecting the functional form of an approximate internal energy. For instance, the most notable departure of $U^{\text{QP(2)}}$ in Eq. (LABEL:QP:U) from the lengthy, but full $U^{(2)}$ of Eq. (74) of Ref. Hirata and Jha, 2020 is the absence of the so-called anomalous-diagram terms,Kohn and Luttinger (1960); Luttinger and Ward (1960) which sum over the set of indexes whose fictitious denominator is zero.Hirata and Jha (2020); Hirata (2021a) They are purposefully neglected in our ansatz because they are found to cause a severe Kohn–Luttinger nonconvergence problem in $\Sigma_{pp}^{(2)}$,Kohn and Luttinger (1960); Luttinger and Ward (1960); Hirata (2021b, 2022) and are undesirable from both theoretical and practical points of view. This issue will be expounded in Appendix A.
When the reference wave function is supplied by a finite-temperature theory, its $\epsilon_{p}^{(0)}$ depends on temperature and the formulas for the thermal self-energies need to be adjusted accordingly. In the thermal HF reference, the second-order thermal self-energy $\Sigma_{pp}^{(2)}$ is found to contain third-order energy-independent diagramsPurvis and Öhrn (1975); Schirmer and Angonoa (1989) among others (not considered further in this study).
Not only does finite-temperature MBPT suffer from the aforementioned Kohn–Luttinger nonconvergence,Hirata (2021b, 2022) but the Feynman–Dyson perturbation series of zero-temperature MBGF is also shown to display even severer divergence for most low- and high-lying states.Hirata et al. (2024) The impact of the latter divergence on thermal quasi-particle theories will be touched upon in Appendix B.
V Thermodynamic functions
Table 1: Grand potential $\Omega$ (in $E_{\text{h}}$) of the hydrogen fluoride molecule (0.9168 Å) in the STO-3G basis set as a function of temperature ($T$).
$T$ / K QP(1)=HF [^1] First-order thermal quasi-particle theory or thermal HF theory. MBPT(2) [^2] Second-order finite-temperature MBPTHirata and Jha (2019, 2020); Hirata (2021a) with zero-temperature HF reference. QP(2) [^3] Second-order thermal quasi-particle theory with zero-temperature HF reference. FCI [^4] Thermal FCI theory.Kou and Hirata (2014)

### v2

@@ Abstract
The widely used thermal Hartree–Fock (HF) theory is generalized to include the effect of electron correlation while maintaining its quasi-independent-particle framework. An electron-correlated internal energy (or grand potential) is postulated in consultation with the second-order finite-temperature many-body perturbation theory (MBPT), which then dictates the corresponding thermal orbital (quasi-particle) energies in such a way that all fundamental thermodynamic relations are obeyed. The associated density matrix is of a one-electron type, whose diagonal elements take the form of the Fermi–Dirac distribution functions, when the grand potential is minimized. The formulas for the entropy and chemical potential are unchanged from those of Fermi–Dirac or thermal HF theory. The theory thus constitutes a finite-temperature extension of the second-order Dyson self-energy of one-particle many-body Green’s function theory and can be viewed as a second-order, diagonal, frequency-independent, thermal inverse Dyson equation. At low temperature, the theory approaches finite-temperature MBPT of the same order, but it may outperform the latter at intermediate temperature by including additional electron-correlation effects through orbital energies. A physical meaning of these thermal orbital energies is proposed (encompassing that of thermal HF orbital energies, which has been elusive) as a finite-temperature version of Janak’s theorem.
@@ I Introduction
Thermal Hartree–Fock (HF) theoryHusimi (1940); Mermin (1963); Gu and Hirata (2024) is a curious ansatz. It uses an auxiliary one-electron Hamiltonian to define its one-electron density matrix, whereas a true density matrix for thermodynamics governed by an interacting Hamiltonian should be much higher ranked.Farid, March, and Theophilou (2000) Its diagonal elements, as variational parameters, ultimately become the Fermi–Dirac distribution functions of the non-interacting Fermi–Dirac theory, when the grand potential is minimized. The state energies defining this grand potential are, however, evaluated with the exact Hamiltonian containing two-electron interactions. Stipulated in this way, the theory does not seem to correspond to a single, consistent grand partition function.
@@
Nevertheless, the thermodynamic functions of this theory are shownArgyres, Kaplan, and Silva (1974); Gu and Hirata (2024) to obey all fundamental thermodynamic relations. They also correctly reduceGu and Hirata (2024) to the zero-temperature HF theory.Szabo and Ostlund (1982); Shavitt and Bartlett (2009) They are widely used in applications to, e.g., energy bands in a solid where their temperature-dependent orbital energies are often invoked, somewhat unquestioningly, in explaining metal-insulator transitions,Hermes and Hirata (2015) etc. This is in spite of the facts that their physical meaning is still obscurePain (2011) and that state energies in quantum mechanics are supposed to be constant of temperature.
@@
In this article, inspired by the immense success of thermal HF theory, we generalize this ansatz to include the effects of electron correlation, while keeping to its quasi-independent-particle framework and grand canonical ensemble. Such a theory, which can provide electron-correlated (i.e., quasi-particle) energy bands of semiconductors at finite temperatures in addition to their thermodynamic functions, may prove even more useful than thermal HF theory or thermal density-functional theory (DFT). Our strategy of postulating what we call thermal quasi-particle theory is as follows:
@@
We first assume that every state consists of non-interacting quasi-particles, each occupying a well-defined orbital with an electron-correlated orbital energy. An internal energy (or grand potential) is postulated by consulting with finite-temperature many-body perturbation theory (MBPT).Hirata and Jha (2019, 2020); Hirata (2021a) Its definition dictates the forms of the thermal quasi-particle energies (i.e., the electron-correlated thermal orbital energies) such that all fundamental thermodynamic relations are obeyed. The corresponding density matrix is then of a one-electron type, whose diagonal elements take the form of the Fermi–Dirac distribution functions, when the grand potential is minimized. The formulas for the entropy and average number of electrons (determining the chemical potential) are unchanged from the corresponding formulas of Fermi–Dirac or thermal HF theory. The thermal quasi-particle energies thus defined are a finite-temperature generalizationMatsubara (1955); Luttinger and Ward (1960); March, Young, and Sampanthar (1967); Kadanoff and Baym (2018) of the Dyson self-energies of one-particle many-body Green’s function (MBGF) theory.Linderberg and Öhrn (1973); Paldus and Čížek (1975); Cederbaum and Domcke (1977); Simons (1977); Öhrn and Born (1981); Jørgensen and Simons (1981); Oddershede (1987); Aryasetiawan and Gunnarsson (1998); Ortiz (1999); Onida, Reining, and Rubio (2002); Ortiz (2013); Hirata et al. (2017, 2024)
@@
Being based on perturbation theory, thermal quasi-particle theory forms a hierarchy of size-consistentHirata (2011) approximations with increasing accuracy and complexity. Fermi–Dirac and thermal HF theories can be viewed, respectively, as the zeroth- and first-order instances of this hierarchy. Here, we introduce a second-order thermal quasi-particle theory, which is the leading order in describing electron correlation. It is based on the second-order grand potential or internal energy of finite-temperature MBPT.Hirata and Jha (2020); Hirata (2021a) We propose the expression for the corresponding second-order thermal self-energies, which obey all thermodynamic relations and ensure the variationality of the grand potential. The theory thus constitutes a second-order, diagonal, frequency-independent, thermal inverse Dyson equation.Hirata et al. (2017) It reduces to the second-order MBPT for internal energy and to the second-order MBGF (in the diagonal and frequency-independent approximation) for orbital energies at zero temperature.
@@
A comparison of thermal quasi-particle theories with finite-temperature MBPT suggests that the former may include higher-order perturbation corrections through correlation-corrected thermal orbital energies and outperform the latter at intermediate temperature. We also propose a physical meaning of these thermal orbital energies in the form of a finite-temperature version of Janak’s theorem.Janak (1978) It encompasses the physical meaning of thermal HF orbital energies, which has been elusive.Pain (2011); Hirata (2021a)
II Thermal quasi-particle theory
II.1 Zeroth order
@@
Fermi–Dirac theory is the zeroth-order thermal quasi-particle theory [thermal QP(0) theory]. It is reviewed briefly to outline the general framework of the hierarchical approximations; it is discussed more fully in Ref. Gu and Hirata, 2024.
Its internal energy $U^{(0)}$ and entropy $S^{(0)}$ at the inverse temperature $\beta=(k_{\text{B}}T)^{-1}$ are given by
@@
where $\bar{N}$ is the average number of electrons that cancel nuclear charges exactly, and the summations are always taken over all spinorbitals.
The $\langle E^{(n)}\rangle$ in Eq. (1) is a zeroth-order thermal average of state energies, i.e.,
@@
where $E_{I}^{(n)}$ is the $n$th-order perturbation correction to the energy of the $I$th Slater-determinant state according to Hirschfelder–Certain degenerate perturbation theory,Hirschfelder and Certain (1974) and $E_{I}^{(0)}=\sum\epsilon_{i}^{(0)}$ ($i$ labels spinorbitals occupied in the $I$th state). Sum-over-states expressions of the thermal averages can be reduced to sum-over-orbitals ones either by combinatorial logicHirata and Jha (2019, 2020) or by normal-ordered second quantization at finite temperature.Hirata (2021a)
@@
$\Omega^{(0)}$$\equiv$ $U^{(0)}-\mu^{(0)}\bar{N}-TS^{(0)}$
$=$ $\sum_{p}\left(\epsilon_{p}^{(0)}-\mu^{(0)}\right)f_{p}^{-}+\frac{1}{\beta}\sum_{p}\left(f_{p}^{-}\ln f_{p}^{-}+f_{p}^{+}\ln f_{p}^{+}\right).$
As explicitly shown in Ref. Gu and Hirata, 2024, they satisfy the thermodynamic relations such as Eq. (II.1) and
@@
which is also explicitly verifiable.Gu and Hirata (2024)
@@
II.2 First order
Thermal HF theoryHusimi (1940); Mermin (1963) constitutes the first-order thermal quasi-particle theory [thermal QP(1) theory]. It is also fully discussed in Ref. Gu and Hirata, 2024. It accommodates the exact Hamiltonian with two-electron interactions within a quasi-independent-particle framework.
Its internal energy $U^{\text{HF}}$ is postulated by an intuitively natural finite-temperature generalization of the zero-temperature HF energy,Hirata (2021a) i.e.,
@@ II.2 First order
$\epsilon_{pq}^{\text{HF}}$ $=$ $h_{pq}+\sum_{r}\langle pr||qr\rangle f_{r}^{-}=\delta_{pq}\epsilon_{p}^{\text{HF}},$ (15)
where $\delta_{pq}$ is Kronecker’s delta.
@@
The last identity underscores that $\Omega^{\text{HF}}$ is minimizedMermin (1963); Gu and Hirata (2024) by the orbitals that diagonalize $\bm{\epsilon}^{\text{HF}}$ [Eq. (15)] and by the one-particle density matrix whose eigenvalues are the Fermi–Dirac distribution functions of Eq. (16).
II.3 Second order
We postulate a second-order thermal quasi-particle theory [thermal QP(2) theory] by its internal energy,
@@ II.3 Second order
The last term $\langle E^{(2)}\rangle$ is the thermal average of second-order Hirschfelder–Certain degenerate perturbation energiesHirschfelder and Certain (1974) and has been reduced in Eqs. (C7) and (C8) of Ref. Hirata and Jha, 2020 to the following sum-over-orbitals formula:
@@
It is a natural finite-temperature generalization of second-order many-body perturbation energy,Bartlett (1981); Shavitt and Bartlett (2009) but is considerably simpler than the full second-order correction to the internal energy $U^{(2)}$ of finite-temperature MBPT [Eq. (74) of Ref. Hirata and Jha, 2020]. The simplification is at least partly justified by the equally simpler treatments of the chemical-potential and entropy contributions to adhere to the quasi-independent-particle framework that the theory adopts.
The first term of $\langle E^{(2)}\rangle$ [Eq. (LABEL:E2)] is the so-called non-HF term (or more precisely, non-thermal-HF term, in this case),Shavitt and Bartlett (2009) which is zero when thermal HF theory is used as the reference (where $F_{pq}=0$). The diagrammatic representationHirata (2021a) of $\langle E^{(2)}\rangle$ is given in Fig. 1.
@@
where $\mu^{\text{QP(2)}}$ is the chemical potential, which is determined by the same electroneutrality condition of Fermi–Dirac theory, which reads
@@
Here, $\Sigma_{pp}^{(2)}$ is a finite-temperature analogue of the second-order Dyson self-energy in the Feynman–Dyson perturbation expansion of MBGF.Hirata et al. (2017) In a zero-temperature reference, where $\epsilon_{p}^{(0)}$ is constant of $T$, it takes the form of
$\Sigma_{pp}^{(2)}$ $=$ $\sum_{q}^{\text{denom.}\neq 0}\frac{F_{qp}F_{pq}}{\epsilon_{p}^{(0)}-\epsilon_{q}^{(0)}}f_{q}^{+}-\sum_{q}^{\text{denom.}\neq 0}\frac{F_{pq}F_{qp}}{\epsilon_{q}^{(0)}-\epsilon_{p}^{(0)}}f_{q}^{-}$ (36)
$+\sum_{q,r}^{\text{denom.}\neq 0}\frac{\langle qp||rp\rangle F_{rq}}{\epsilon_{r}^{(0)}-\epsilon_{q}^{(0)}}f_{r}^{-}f_{q}^{+}$
$+\sum_{q,r}^{\text{denom.}\neq 0}\frac{F_{qr}\langle rp||qp\rangle}{\epsilon_{r}^{(0)}-\epsilon_{q}^{(0)}}f_{r}^{-}f_{q}^{+}$
@@
which is obtained by demanding all thermodynamic relations be obeyed (see below for a derivation). A diagrammatic representation of $\Sigma_{pp}^{(2)}$ is given in Fig. 2.
@@
When a zero-temperature HF reference is used, in the $T=0$ limit, the first four terms vanish because $F_{pq}=0$, leaving only the last two terms as the zero-temperature second-order self-energy in the diagonal, frequency-independent approximation. Their diagrams (the last two in Fig. 2) are, therefore, isomorphic to the two second-order self-energy diagrams (e.g., Fig. 4 of Ref. Hirata et al., 2017) of zero-temperature MBGF, although their scope is different (the former are for $T\geq 0$, while the latter are only for $T=0$).
Figure 2: Second-order thermal Dyson self-energy diagrams in the zero-temperature reference [Eq. (36)], which are obtained by deleting an edge in each of the internal energy diagrams in Fig. 1. See also the caption of Fig. 1 as well as Refs. Hirata, 2021a; Hirata et al., 2017.
@@
but not the explicit derivative $\partial f_{p}^{-}/\partial\mu^{\text{QP(2)}}$, which is a solution of a system of linear equations and cannot be written in a closed form.Gu and Hirata (2024) This justifies the condition for $\mu^{\text{QP(2)}}$ [Eq. (32)].
@@
which can be verified by using Eqs. (LABEL:E2) and (36). In fact, the form of $\Sigma_{pp}^{(2)}$ is determined such that the above equation is satisfied in the first place. Generally,
@@
which ensures that Eq. (41) and all similar relations hold.
Equation (42) is a finite-temperature generalizationMatsubara (1955); Luttinger and Ward (1960) of the $n$th-order self-energy of Feynman–Dyson MBGF in the diagonal, frequency-independent approximation.Hirata et al. (2017, 2015) Owing to this diagonal construction, the second- and higher-order thermal quasi-particle theories do not involve rotation of orbitals, unlike thermal HF theoryMermin (1963); Gu and Hirata (2024) or Feynman–Dyson MBGF without the diagonal approximation,Hirata et al. (2017) which define the HF or Dyson orbitalsOrtiz (2020) as those that bring the Fock or self-energy matrix into a diagonal form. Despite the absence of orbital rotation, $\Omega^{\text{QP(2)}}$ is still minimized by varying $f_{p}^{-}$ (see below).
Diagrammatically, the differentiation with respect to $f_{p}^{-}$ corresponds to opening a closed, internal-energy diagram by deleting an edge (whose mathematical interpretation is $f_{p}^{\pm}$).Hirata et al. (2017) A physical meaning of thermal self-energies or thermal quasi-particle energies, which are temperature dependent, is discussed in Sec. III.2.
@@
$-k_{\text{B}}\beta^{2}\sum_{p}\Sigma_{pp}^{(2)}\frac{\partial f_{p}^{-}}{\partial\beta}+k_{\text{B}}\beta^{2}\frac{\partial\langle E^{(2)}\rangle}{\partial\beta}$
@@
where we used Eqs. (39), (40), and (42), but not the explicit derivative $\partial f_{p}^{-}/\partial\beta$.Gu and Hirata (2024) This justifies the entropy formula [Eq. (29)].
@@
$\frac{\partial\Omega^{\text{QP(2)}}}{\partial f_{p}^{-}}$ $=$ $\left(\epsilon_{p}^{\text{HF}}-\mu^{\text{QP(2)}}\right)+\sum_{q}\frac{\partial\epsilon_{q}^{\text{HF}}}{\partial f_{p}^{-}}f_{q}^{-}-\sum_{q}\langle pq||pq\rangle f_{q}^{-}$
$+\frac{1}{\beta}\ln\frac{f_{p}^{-}}{f_{p}^{+}}+\frac{\partial\langle E^{(2)}\rangle}{\partial f_{p}^{-}}$
@@
where Eq. (42) was used. This proves that the $f_{p}^{-}$ functions that minimize $\Omega^{\text{QP(2)}}$ take the forms of the Fermi–Dirac distribution functions, justifying Eqs. (33)–(35).
Insofar as Eq. (42) is satisfied (and thus all thermodynamic relations are obeyed and the variationality is ensured), there is considerable latitude in selecting the functional form of an approximate internal energy. For instance, the most notable departure of $U^{\text{QP(2)}}$ in Eq. (LABEL:QP:U) from the lengthy, but full $U^{(2)}$ of Eq. (74) of Ref. Hirata and Jha, 2020 is the absence of the so-called anomalous-diagram terms,Kohn and Luttinger (1960); Luttinger and Ward (1960) which sum over the set of indexes whose fictitious denominator is zero.Hirata and Jha (2020); Hirata (2021a) They are purposefully neglected in our ansatz because they are found to cause a severe Kohn–Luttinger nonconvergence problemKohn and Luttinger (1960); Luttinger and Ward (1960); Hirata (2021b, 2022) in $\Sigma_{pp}^{(2)}$, and are undesirable. This issue will be expounded in Appendix A.
When the reference wave function is supplied by a finite-temperature theory, its $\epsilon_{p}^{(0)}$ depends on temperature and the formulas for the thermal self-energies need to be adjusted accordingly. In the thermal HF reference, the second-order thermal self-energy $\Sigma_{pp}^{(2)}$ is found to contain third-order energy-independent terms.Purvis and Öhrn (1975); Schirmer and Angonoa (1989) Although they seem no more expensive to evaluate than the rest, we will not consider this reference any further in this initial study.
Not only is finite-temperature MBPT plagued by the Kohn–Luttinger nonconvergence,Hirata (2021b, 2022) but the Feynman–Dyson perturbation series of zero-temperature MBGF is also shown to display even severer divergence for most low- and high-lying states.Hirata et al. (2024) Potential impact of the latter divergence on thermal quasi-particle theories will be contemplated in Appendix B.
III Numerical results
III.1 Thermodynamic functions
Table 1: Grand potential $\Omega$ (in $E_{\text{h}}$) of an ideal gas of the identical hydrogen fluoride molecules (0.9168 Å) in the STO-3G basis set as a function of temperature ($T$).
$T$ / K HF [^1] Thermal HF theory. MBPT(2) [^2] Second-order finite-temperature MBPTHirata and Jha (2019, 2020); Hirata (2021a) with zero-temperature HF reference. QP(2) [^3] Second-order thermal quasi-particle theory with zero-temperature HF reference. FCI [^4] Thermal FCI theory.Kou and Hirata (2014)